OpenStudy (shamil98):

\[\ \int_{}^{} \cos^4 x -\sin^4 x ~~dx\] i'm supposed to use this identity along with u-sub \[\ \cos 2x =cos^2 x- sin^2 x \]

OpenStudy (shamil98):

er one sec latex messed up

OpenStudy (shamil98):

nvm shows now

OpenStudy (anonymous):

rewrite the function in terms of cos^2x-sin^2x

OpenStudy (anonymous):

\[\cos ^{2}x-\sin ^{2}x\]

OpenStudy (shamil98):

wouldn't work would it? since it's a -sin^2x \[\int (cos^2x - sin^2x)^2~~dx\]

OpenStudy (shamil98):

or is the negative sign just not important

OpenStudy (anonymous):

you have to add \[\cos ^{2}xsin ^{2}x\]

OpenStudy (anonymous):

I'm gonna let someone else take this one

OpenStudy (loser66):

I have an idea : \((cos^2)-(sin^2)^2 =(cos^2+sin^2)(cos^2-sin^2) = cos^2-sin^2= cos(2x)\)

OpenStudy (loser66):

the first term is \((cos^2)^2\)

OpenStudy (shamil98):

ah

OpenStudy (shamil98):

I thought of that as well

OpenStudy (shamil98):

\[\ \int \cos^2 2x~~ dx \] it ends up like this right?

OpenStudy (loser66):

nope, just cos (2x)

OpenStudy (shamil98):

i mean since \[\ cos^2 x - sin^2x = cos 2x \] then \[\ cos^4x - sin^4x = cos^2x\]

OpenStudy (loser66):

I mean \(\ \int \cos( 2x)~~ dx\)

OpenStudy (shamil98):

oops typo \[\ cos^4 x - sin^4 x = cos^2 2x\]

OpenStudy (shamil98):

does that work?

OpenStudy (loser66):

hahahah, how???

OpenStudy (loser66):

Ok, let me write the whole thing out

OpenStudy (shamil98):

kk

OpenStudy (noelgreco):

Factor the original integrand as a difference of squares. One factor is = 1 (Trig ident) and the other is cos(2x)

OpenStudy (loser66):

\((cos^2(x))^2-(sin^2(x))^2= (cos^2(x)+sin^2(x)(cos^2(x)-sin^2(x))\\=1*(cos^2(x)-sin^2(x) = cos(2x)\)\)

OpenStudy (loser66):

like \(a^4 -b^4 = (a^2)^2 -(b^2)^2=(a^2+b^2)(a^2-b^2)\)

OpenStudy (loser66):

your a is cos, your b is sin, hence the first term a^2 +b^2 =1

OpenStudy (noelgreco):

Yes

OpenStudy (shamil98):

Oh... it makes sense now lol thnx

OpenStudy (loser66):

hihihi.... students have that phase, me too, hihihi